OpenStudy (anonymous):

-3(2x-7)=-4+37-2(x+8) need help solving and what are the solutions?

OpenStudy (lilg132):

from what i can tell x = 1

OpenStudy (lilg132):

so first you expand the brackets: -3(2x - 7) = -4 + 37 - 2(x + 8) Lets expand the left side before the equals sign first -3(2x - 7) Will be expanded to look like this -6x + 21 Now lets expand the right side after the equals sign -4 + 37 - 2(x + 8) Will be expanded to look like this -4 + 37 - 2x - 16 Now what we have is : -6x + 21 = -4 + 37 - 2x - 16 In order to solve x we need to bring all the x's to one side and all the whole numbers to the other side.(remember when moving to the other side it will be the opposite of what it was e.g. moving -4 to the other side will be +4) so what we have is : 4 - 37 + 16 + 21 = 6x - 2x Now we just add and subtract as written so: 4 - 37 + 16 + 21 = 4 6x - 2x = 4x so now we have: 4 = 4x In order to work out what x is we have to divide by whatever x is which is 4 so we divide both sides by 4 4 / 4 = 4x / 4 which equals to : 1 = x Please dont forget to give me a medal or good answer :D

OpenStudy (anonymous):

I will definitley give you a medal, i was confused on how to do this, I am new to algebra..Thank you bunches lilg..can I add you as friend for help?

OpenStudy (lilg132):

yes no problem

OpenStudy (anonymous):

is the final answer is x=1 and not 1=x? just varifying...

OpenStudy (anonymous):

i got it..thanks again lilg 132

OpenStudy (lilg132):

x = 1 is the correct answer

OpenStudy (anonymous):

:)